# On automatic boundedness of operators from ordered to topological vector spaces

### 16 Apr 2026 v3

@@ 1. $i)$
    If $X_{+}$ is normal and $Y$ is a NS, then $\text{\bf L}_{rnc}(X,Y)=\text{\bf L}_{onb}(X,Y)=\text{\bf L}_{b}(X,Y)$ [12, Corollary 2.12].
@@ 5. $v)$
    If $X_{+}$ is normal and $Y=(Y,\tau)$ is a Banach space with the weak topology, then ${\cal L}_{otc}^{\sigma}(X,Y)\subseteq{\cal L}_{onb}(X,Y)$ [12, Proposition 2.16].
@@ 6. $vi)$
$\text{\bf L}_{otb}({\cal F},Y)\subseteq\text{\bf L}_{c}({\cal F},Y)$, whenever $({\cal F},\tau)$ is an ordered Fréchet space with a $\tau$-closed generating cone and $(Y,\xi)$ is a topological vector space.
@@
Suppose, on the contrary, ${\cal T}\in\text{\bf L}_{otb}({\cal F},Y)\setminus\text{\bf L}_{c}({\cal F},Y)$. For each $n\in\mathbb{N}$, denote
@@
Clearly, $\{B_{\cal F}(0,2^{-2n}):n\in\mathbb{N}\}$ is a $\tau$-neighborhood base of zero consisting of $\tau$-closed sets and $B_{\cal F}(0,2^{-2(n+1)})+B_{\cal F}(0,2^{-2(n+1)})\subseteq B_{\cal F}(0,2^{-2n})$ holds for each $n$. Furthermore, each $B_{\cal F}(0,2^{-2n})$ is circled. To see this, it suffices to show that $p(x,0)\leq 2^{-2n}$ implies $p(tx,0)\leq 2^{-2n}$ whenever $|t|\leq 1$. So, let $p(x,0)\leq 2^{-2n}$ and $|t|\leq 1$. Since $\frac{|t|a}{1+|t|a}\leq\frac{a}{1+a}$ for every $a\in\mathbb{R}_{+}$, we obtain $p(tx,0)=\sum\limits_{k=1}^{\infty}\frac{2^{-k}p_{k}(tx)}{1+p_{k}(tx)}=\sum\limits_{k=1}^{\infty}\frac{2^{-k}|t|p_{k}(x)}{1+|t|p_{k}(x)}\leq\sum\limits_{k=1}^{\infty}\frac{2^{-k}p_{k}(x)}{1+p_{k}(x)}=p(x,0)\leq 2^{-2n}$. By the Ando theorem [4, Theorem 2.10], $\{V_{n}:n\in\mathbb{N}\}$ is a $\tau$-neighborhood base of zero consisting of (not necessarily open) circled subsets satisfying $V_{n+1}+V_{n+1}\subseteq V_{n}$ for all $n$.
@@
Every order bounded operator from an ordered Fréchet space with a closed generating cone to an ordered locally convex space with a normal cone is continuous.
@@ 3 On continuity of the Levi and the Lebesgue operators
We shall use the following versions of $\tau$-Levi and $\tau$-Lebesgue operators (see also [5, 11, 19, 21] in the vector lattice-setting).
@@ 1. $a)$
    quasi ($\sigma$-)Levi if $T$ takes $\tau$-bounded increasing nets ((sequences)) of $X$ to o-Cauchy ones (briefly, $T\in{\cal L}_{\text{\rm qLevi}}(X,Y)$ (resp., $T\in{\cal L}^{\sigma}_{\text{\rm qLevi}}(X,Y))$.
@@ 1. $b)$
    $(\sigma$-))Lebesgue if $Tx_{\alpha}\operatorname{\xrightarrow[]{\xi}}0$ for every ((sequence)) net $x_{\alpha}\downarrow 0$ in $X$ [5, Definition 1.1]. The set of such operators is denoted by ${\cal L}_{Leb}(X,Y)$ (resp., ${\cal L}^{\sigma}_{Leb}(X,Y)$).
@@ 2. $c)$
    $(\sigma$-))w-Lebesgue if $Tx_{\alpha}\operatorname{\xrightarrow[]{\sigma(Y,Y^{\prime})}}0$ for every ((sequence)) net $x_{\alpha}\downarrow 0$ in $X$ [5, Definition 1.1]. The set of such operators is denoted by ${\cal L}_{wLeb}(X,Y)$ (resp., ${\cal L}^{\sigma}_{wLeb}(X,Y)$).
@@ 3. $d)$
    $(\sigma$-))order-to-weak continuous if $Tx_{\alpha}\operatorname{\xrightarrow[]{\sigma(Y,Y^{\prime})}}0$ for every ((sequence)) net $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$. The set of such operators is denoted by ${\cal L}_{owc}(X,Y)$ (resp., ${\cal L}^{\sigma}_{owc}(X,Y)$) (cf., Definition 2.1d)d)).
We prove that, under rather mild conditions, quasi $\sigma$-Levi operators are continuous. This result generalizes [11, Theorem 3.1].
@@
Every quasi $\sigma$-Levi operator from a Banach lattice to a normed lattice is continuous.
Proof.
Let $T:E\to F$ be a quasi $\sigma$-Levi operator from a BL $E$ to an NL $F$. On the way to a contradiction, suppose $T$ is not continuous. Then, there exists a sequence $(x_{n})$ in $B_{X}$ such that $\|Tx_{n}\|\to\infty$. Without loss of generality, suppose $(x_{n})\subset(B_{X})^{+}$ and $\|Tx_{n}\|\geq n2^{n}$ for every $n$. Set $u=\sum\limits_{k=1}^{\infty}2^{-k}x_{k}$ and $u_{n}=\sum\limits_{k=1}^{n}2^{-k}x_{k}$ for each $n\in\mathbb{N}$. Then $0\leq u_{n}\uparrow$ and $\|u_{n}\|\leq 1$ for all $n$. Since $T\in\text{\rm L}_{\text{\rm qLevi}}(E,F)$, $(Tu_{n})$ is an o-Cauchy sequence in $F$, and hence the sequence $(Tu_{n})$ is order bounded. Thus, $(Tu_{n})$ is also norm bounded. It is absurd because $\|Tu_{n+1}-Tu_{n}\|=\|T2^{-(n+1)}x_{n+1}\|\geq n+1$ for all $n$. ∎
Theorem 3.3.
Every quasi $\sigma$-Levi operator from an ordered Fréchet space with a closed generating cone to an ordered normed space with a normal cone is continuous.
@@
It follows from $T\in{\cal L}^{\sigma}_{\text{\rm qLevi}}({\cal F},Y)$ that $(Tu_{i})$ is o-Cauchy in $Y$, and hence is order bounded. Since $Y_{+}$ is normal, the sequence $(Tu_{i})$ is also norm bounded. It is absurd because
$\|Tu_{i+1}-Tu_{i}\|=\|Tx_{i+1}\|=\|T(2^{i+1}z_{i+1})\|\geq 2^{i+1}M$
for all $i\in\mathbb{N}$. ∎
@@
Let $T$ be a positive operator from an OVS $X$ to an ordered TVS $(Y,\tau)$ with a normal cone. Then
@@
On the way to a contradiction, suppose $T\in{\cal L}_{Leb}(X,Y)\setminus{\cal L}_{otc}(X,Y)$. Then, there exist a net $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$ and a neighborhood $U\in\tau(0)$ such that, for every $\alpha$, there exists $\alpha^{\prime}\geq\alpha$ with $Tx_{\alpha^{\prime}}\notin U$. Since $Y_{+}$ is normal, we may suppose that $U$ is circled and full by [4, Corollary 2.21].
Pick a net $g_{\beta}\downarrow 0$ in $X$ such that, for each $\beta$, there exists $\alpha_{\beta}$ with $x_{\alpha}\in[-g_{\beta},g_{\beta}]$ for $\alpha\geq\alpha_{\beta}$. As $T$ is Lebesgue, $Tg_{\beta}\operatorname{\xrightarrow[]{\tau}}0$. Since $T\geq 0$, $Tg_{\beta}\downarrow\geq 0$. Suppose $Tg_{\beta}\geq f\geq 0$ for some $f\in Y$. Since $\tau$ is a Hausdorff topology, it follows from [4, Theorem 2.23] that $f=0$, and hence $Tg_{\beta}\downarrow 0$. Since $T\geq 0$, we obtain $Tx_{\alpha}\in[-Tg_{\beta},Tg_{\beta}]$ for $\alpha\geq\alpha_{\beta}$. Since $Tg_{\beta}\operatorname{\xrightarrow[]{\tau}}0$, there exists $\beta_{U}$ such that $Tg_{\beta}\in U$ for all $\beta\geq\beta_{U}$. Since $U$ is circled and full, $Tg_{\beta_{U}},Tg_{\beta_{U}}]\subseteq U$ for all $\beta\geq\beta_{U}$, and hence $Tx_{\alpha}\in U$ for all $\alpha\geq\alpha_{\beta_{U}}$. It is absurd because $Tx_{\alpha^{\prime}_{\beta_{U}}}\notin U$. Thus, $T\in{\cal L}_{otc}(X,Y)$. ∎
@@
Let $X$ be an OVS and $Y$ be a normal ONS. Then $({\cal L}_{Leb})_{+}(X,Y)=({\cal L}_{onc})_{+}(X,Y)$ and $({\cal L}^{\sigma}_{Leb})_{+}(X,Y)=({\cal L}^{\sigma}_{onc})_{+}(X,Y)$.
@@
As the $\sigma$-Lebesgue case is similar, we consider only the first equivalence. Suppose $T\in{\cal L}_{wLeb}(X,Y)$ and take $x_{\alpha}\downarrow 0$ in $X$. Since $T\geq 0$, we have $Tx_{\alpha}\downarrow\geq 0$ and, as $T$ is w-Lebesgue, $Tx_{\alpha}\operatorname{\xrightarrow[]{\sigma(Y,Y^{\prime})}}0$. By [4, Lemma 2.28], we conclude $Tx_{\alpha}\operatorname{\xrightarrow[]{\tau}}0$. Thus, $T\in{\cal L}_{Leb}(X,Y)$. ∎
@@
Question 3.8.
Are the conclusions of Corollaries 3.5 and 3.7 hold true for arbitrary (not necessarily positive) operators?
@@
Let $X$ be an OBS with a closed generating normal cone and let $Y$ be an NS. Then ${\cal L}_{onb}(X,Y)={\cal L}_{b}(X,Y)$.
@@
The inclusion ${\cal L}_{onb}(X,Y)\subseteq{\cal L}_{b}(X,Y)$ follows from [15, Corollary 2.9]. The reverse inclusion hods due to normality of $X$. ∎
By [15, Theorem 2.8], every order-to-norm bounded operator from an OBS with closed generating cone to an NS is bounded. Accordingly to [15, Theorem 2.4], each order bounded operator from an OVS to an OVS with generating cones is ru-continuous. Combining these two results and keeping in mind that order intervals in an OBS with normal cone are bounded, we obtain the following automatic boundedness conditions of order bounded operators:
  1. (*)
    ${\cal L}_{ob}(X,Y)\subseteq{\cal L}_{b}(X,Y)$, whenever $X$ and $Y$ are OBSs with generating cones such that $X_{+}$ is closed and $Y_{+}$ is normal.
Another case of automatic boundedness of operators concerns the order continuity. Namely, as every order continuous operator from an Archimedean OVS with generating cone to an OVS is order bounded [15, Corollary 2.2], we have:
  1. (**)
    ${\cal L}_{oc}(X,Y)\subseteq{\cal L}_{b}(X,Y)$, whenever $X$ and $Y$ are OBSs with generating cones such that $X_{+}$ is closed and $Y_{+}$ is normal.
Condition $X_{+}-X_{+}=X$ is essential in both (*) and (**) since each $T:X\to X$ is order bounded and order continuous whenever $X_{+}=\{0\}$. An inspection of [15, Example 2.12 b)] tells us that each infinite dimensional Banach space $X$ has a non-closed generating Archimedean cone so that every operator $T:X\to X$ is order continuous, and hence order bounded.
It is worth nothing that the order-to-norm and order-to-weak continuous operators on vector lattices were defined in [21] without specifying the class of (just linear or continuous linear) operators. Automatic boundedness of such operators was investigated in recent paper [22].
We are going to establish some automatic boundedness conditions for order-to-weak continuous operators on OVSs. More precisely, we shall prove Theorem 3.12 which tells us that every $\sigma$-w-Lebesgue operator from an OBS with a closed generating normal cone to an NS is bounded. We begin with conditions under which $\sigma$-w-Lebesgue operators are order-to-norm bounded.
@@
Let $X$ be a normal OBS and $Y$ be an NS. Then ${\cal L}^{\sigma}_{wLeb}(X,Y)\subseteq{\cal L}_{onb}(X,Y)$. In particular, ${\cal L}^{\sigma}_{owc}(X,Y)\subseteq{\cal L}_{onb}(X,Y)$.
@@
Let $T:X\to Y$ be $\sigma$-w-Lebesgue. If $T$ is not order-to-norm bounded then $T[0,u]$ is not bounded for some $u\in X_{+}$. Since $X$ is normal then $[0,u]$ is bounded, say $\sup\limits_{x\in[0,u]}\|x\|\leq M$. Pick a sequence $(u_{n})$ in $[0,u]$ with $\|Tu_{n}\|\geq n2^{n}$, and set $y_{n}:=\|\cdot\|$-$\sum\limits_{k=n}^{\infty}2^{-k}u_{k}$ for $n\in\mathbb{N}$. Then $y_{n}\downarrow\geq 0$. Let $0\leq y_{0}\leq y_{n}$ for all $n\in\mathbb{N}$. Since $0\leq y_{0}\leq 2^{1-n}u$ and $\|2^{1-n}u\|\leq M2^{1-n}\to 0$ then $y_{0}=0$ by [4, Theorem 2.23]. Thus, $y_{n}\downarrow 0$. It follows from $T\in{\cal L}^{\sigma}_{wLeb}(X,Y)$ that $(Ty_{n})$ is w-null, and hence is norm bounded, that is absurd because $\|Ty_{n+1}-Ty_{n}\|=\|T(2^{-n}u_{n})\|\geq n$ for all $n\in\mathbb{N}$. ∎
Combining Lemmas 3.9, 3.10, and 3.11, we obtain.
@@
Let $X$ be an ordered Banach space with a closed generating normal cone and let $Y$ be a normed space. Then ${\cal L}^{\sigma}_{wLeb}(X,Y)\subseteq{\cal L}_{b}(X,Y)$. In particular, ${\cal L}^{\sigma}_{ows}(X,Y)\subseteq{\cal L}_{b}(X,Y)$.
@@
(cf. [5, Proposition 3.2], [22, Lemma 2.3]) Each $\sigma$-w-Lebesgue, and hence each $\sigma$-order-to-weak continuous, operator from a Banach lattice to an NS is bounded.
The next result is a straightforward consequence of Lemma 3.9 and Theorem 3.12.
Corollary 3.14.
Let $X$ be an OBS with a closed generating normal cone and let $Y$ be an NS. Then ${\cal L}_{Leb}(X,Y)$, ${\cal L}^{\sigma}_{Leb}(X,Y)$, ${\cal L}_{wLeb}(X,Y)$, ${\cal L}^{\sigma}_{wLeb}(X,Y)$, ${\cal L}_{onc}(X,Y)$, ${\cal L}^{\sigma}_{onc}(X,Y)$, ${\cal L}_{owc}(X,Y)$, and ${\cal L}^{\sigma}_{owc}(X,Y)$ are norm closed subspaces of ${\cal L}_{b}(X,Y)$.
Positive order-to-norm continuous order-to-norm bounded operators between normed lattices are not necessary bounded (see [5, Remark 1.1. c)]). From the other hand, even a positive $\sigma$-Lebesgue rank one contraction on an AM-space need not to be w-Lebesgue [5, Example 3.1].
@@
The topology in an ordered TVS $(X,\tau)$ (the norm in an ONS $X$) is
  1. $a)$
    order continuous if $x_{\alpha}\operatorname{\xrightarrow[]{\tau}}0$ (resp., $\|x_{\alpha}\|\to 0$) for every net $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$.
  2. $b)$
    $\sigma$-order continuous if $x_{n}\operatorname{\xrightarrow[]{\tau}}0$ (resp., $\|x_{n}\|\to 0$) for every sequence $x_{n}\operatorname{\xrightarrow[]{o}}0$ in $X$.
@@ 2. $b)$
Let $X$ be an OBS with a closed normal cone. The following assertions hold.
  1. $i)$
    If the norm in $X$ is order continuous then $x_{\alpha}\operatorname{\xrightarrow[]{o}}0\Longleftrightarrow x_{\alpha}\operatorname{\xrightarrow[]{ru}}0$ in $X$.
  2. ii)ii)
    If the norm in $X$ is $\sigma$-order continuous then $x_{n}\operatorname{\xrightarrow[]{o}}0\Longleftrightarrow x_{n}\operatorname{\xrightarrow[]{ru}}0$ in $X$.

### 20 Apr 2026 v4

@@ 1. $i)$
    If $X_{+}$ is normal and $Y$ is a NS, then $\text{\bf L}_{rtc}(X,Y)=\text{\bf L}_{otb}(X,Y)=\text{\bf L}_{b}(X,Y)$ [12, Corollary 2.12].
@@ 5. $v)$
    If $X_{+}$ is normal and $Y=(Y,\tau)$ is a Banach space with the weak topology, then ${\cal L}_{otc}^{\sigma}(X,Y)\subseteq{\cal L}_{otb}(X,Y)$ [12, Proposition 2.16].
@@ 6. $vi)$
$\text{\bf L}_{otb}({\cal F},Y)\subseteq\text{\bf L}_{tc}({\cal F},Y)$, whenever $({\cal F},\tau)$ is an ordered Fréchet space with a $\tau$-closed generating cone and $(Y,\xi)$ is a topological vector space.
@@
Suppose, on the contrary, ${\cal T}\in\text{\bf L}_{otb}({\cal F},Y)\setminus\text{\bf L}_{tc}({\cal F},Y)$. For each $n\in\mathbb{N}$, denote
@@
Clearly, $\{B_{\cal F}(0,2^{-2n}):n\in\mathbb{N}\}$ is a $\tau$-neighborhood base of zero consisting of $\tau$-closed sets and $B_{\cal F}(0,2^{-2(n+1)})+B_{\cal F}(0,2^{-2(n+1)})\subseteq B_{\cal F}(0,2^{-2n})$ holds for each $n$. Furthermore, each $B_{\cal F}(0,2^{-2n})$ is circled. To see this, it suffices to show that $p(x,0)\leq 2^{-2n}$ implies $p(tx,0)\leq 2^{-2n}$ whenever $|t|\leq 1$. So, let $p(x,0)\leq 2^{-2n}$ and $|t|\leq 1$. Since $\frac{|t|a}{1+|t|a}\leq\frac{a}{1+a}$ for every $a\in\mathbb{R}_{+}$, we obtain $p(tx,0)=\sum\limits_{k=1}^{\infty}\frac{2^{-k}p_{k}(tx)}{1+p_{k}(tx)}=\sum\limits_{k=1}^{\infty}\frac{2^{-k}|t|p_{k}(x)}{1+|t|p_{k}(x)}\leq\sum\limits_{k=1}^{\infty}\frac{2^{-k}p_{k}(x)}{1+p_{k}(x)}=p(x,0)\leq 2^{-2n}$. By the Ando theorem [4, Theorem 2.10], $\{V_{n}:n\in\mathbb{N}\}$ is a $\tau$-neighborhood base of zero consisting of circled sets satisfying $V_{n+1}+V_{n+1}\subseteq V_{n}$ for all $n$.
@@
Every order bounded operator from an ordered Fréchet space with a closed generating cone to an ordered topological vector space with a normal cone is continuous.
@@ 3 On continuity of the Levi and the Lebesgue operators
In this section, we study $\tau$-Levi and $\tau$-Lebesgue operators [5, 19, 12] in the more general OVS-setting. We prove Theorem 3.2 saying that a quasi $\sigma$-Levi operator from an ordered Fréchet space with a closed generating cone to an ordered topological vector space with a normal cone is continuous. This theorem is a far going generalization of [11, Theorem 3.1], where the continuity is established for a quasi $\sigma$-Levi operator from a Banach lattice to a normed lattice. We mention also: Lemma 3.11 saying that each $\sigma$-w-Lebesgue operator is order-to-norm bounded whenever an OBS in the domain has a normal cone; Theorem 3.12 saying that, if the domain has closed generating normal cone, then every $\sigma$-w-Lebesgue operator is bounded; and Theorem 3.3 providing conditions for ($\sigma$-)order to (weak) norm continuity of ($\sigma$-)(w-)Lebesgue operators. We shall use below the following versions of $\tau$-Levi and $\tau$-Lebesgue operators (see [5, 11, 19, 21] in the vector lattice-setting).
@@ 1. $a)$
    quasi $\sigma$-Levi if $T$ takes $\tau$-bounded increasing sequences of $X$ to o-Cauchy ones.
@@ 1. $b)$
    $(\sigma$-))Lebesgue if $Tx_{\alpha}{\xrightarrow[]{\tau}}0$ for every ((sequence)) net $x_{\alpha}\downarrow 0$ in $X$. The set of such operators is denoted by ${\cal L}_{Leb}(X,Y)$ (resp., ${\cal L}^{\sigma}_{Leb}(X,Y)$).
@@ 2. $c)$
    $(\sigma$-))w-Lebesgue if $Tx_{\alpha}{\xrightarrow[]{\sigma(Y,Y^{\prime})}}0$ for every ((sequence)) net $x_{\alpha}\downarrow 0$ in $X$. The set of such operators is denoted by ${\cal L}_{wLeb}(X,Y)$ (resp., ${\cal L}^{\sigma}_{wLeb}(X,Y)$).
@@ 3. $d)$
    $(\sigma$-))order-to-weak continuous if $Tx_{\alpha}{\xrightarrow[]{\sigma(Y,Y^{\prime})}}0$ for every ((sequence)) net $x_{\alpha}{\xrightarrow[]{o}}0$ in $X$. The set of such operators is denoted by ${\cal L}_{owc}(X,Y)$ (resp., ${\cal L}^{\sigma}_{owc}(X,Y)$) (cf., Definition 2.2d)d)).
First, we prove that under rather mild conditions, a quasi $\sigma$-Levi operator is continuous. This generalizes [11, Theorem 3.1] which states that every quasi $\sigma$-Levi operator from a Banach lattice to a normed lattice is continuous.
@@
Every quasi $\sigma$-Levi operator from an ordered Fréchet space with a closed generating cone to an ordered topological vector space with a normal cone is continuous.
@@
As $T$ is a quasi $\sigma$-Levi operator, the sequence $(Tu_{i})$ is o-Cauchy in $Y$, and hence is order bounded. Since $Y_{+}$ is normal, the sequence $(Tu_{i})$ is topologically bounded by [4, Lemma 2.22]. It is absurd because $Tu_{i+1}-Tu_{i}=Tx_{i+1}=2^{i+1}Tz_{i+1}\notin 2^{i+1}W$ for all $i\in\mathbb{N}$. ∎
@@
Let $T$ be a positive operator from an ordered vector space $X$ to an ordered topological vector space $(Y,\tau)$ with a normal cone. Then
@@
On the way to a contradiction, suppose $T\in{\cal L}_{Leb}(X,Y)\setminus{\cal L}_{otc}(X,Y)$. Then, there exist a net $x_{\alpha}{\xrightarrow[]{o}}0$ in $X$ and a neighborhood $U\in\tau(0)$ such that, for every $\alpha$, there exists $\alpha^{\prime}\geq\alpha$ with $Tx_{\alpha^{\prime}}\notin U$. Since $Y_{+}$ is normal, we may suppose that $U$ is circled and full by [4, Corollary 2.21].
Pick a net $g_{\beta}\downarrow 0$ in $X$ such that, for each $\beta$, there exists $\alpha_{\beta}$ with $x_{\alpha}\in[-g_{\beta},g_{\beta}]$ for $\alpha\geq\alpha_{\beta}$. As $T$ is Lebesgue, $Tg_{\beta}{\xrightarrow[]{\tau}}0$. Since $T\geq 0$, $Tg_{\beta}\downarrow\geq 0$. Suppose $Tg_{\beta}\geq f\geq 0$ for some $f\in Y$. Since $\tau$ is a Hausdorff topology, it follows from [4, Theorem 2.23] that $f=0$, and hence $Tg_{\beta}\downarrow 0$. Since $T\geq 0$, we obtain $Tx_{\alpha}\in[-Tg_{\beta},Tg_{\beta}]$ for $\alpha\geq\alpha_{\beta}$. Since $Tg_{\beta}{\xrightarrow[]{\tau}}0$, there exists $\beta_{U}$ such that $Tg_{\beta}\in U$ for all $\beta\geq\beta_{U}$. Since $U$ is circled and full, $Tg_{\beta_{U}},Tg_{\beta_{U}}]\subseteq U$ for all $\beta\geq\beta_{U}$, and hence $Tx_{\alpha}\in U$ for all $\alpha\geq\alpha_{\beta_{U}}$. It is absurd because $Tx_{\alpha^{\prime}_{\beta_{U}}}\notin U$. Thus, $T\in{\cal L}_{otc}(X,Y)$. ∎
@@
Let $X$ be an OVS and $Y$ be a normal ONS. Then $({\cal L}_{Leb})_{+}(X,Y)=({\cal L}_{otc})_{+}(X,Y)$ and $({\cal L}^{\sigma}_{Leb})_{+}(X,Y)=({\cal L}^{\sigma}_{otc})_{+}(X,Y)$.
@@
As the $\sigma$-Lebesgue case is similar, we consider only the first equivalence. Suppose $T\in{\cal L}_{wLeb}(X,Y)$ and take $x_{\alpha}\downarrow 0$ in $X$. Since $T\geq 0$, we have $Tx_{\alpha}\downarrow\geq 0$ and, as $T$ is w-Lebesgue, $Tx_{\alpha}{\xrightarrow[]{\sigma(Y,Y^{\prime})}}0$. By [4, Lemma 2.28], we conclude $Tx_{\alpha}{\xrightarrow[]{\tau}}0$. Thus, $T\in{\cal L}_{Leb}(X,Y)$. ∎
@@
Question 3.7.
Are the conclusions of Corollaries 3.4 and 3.6 hold true for arbitrary (not necessarily positive) operators?
@@
Let $X$ be an OBS with a closed generating normal cone and let $Y$ be an NS. Then, ${\cal L}_{otb}(X,Y)={\cal L}_{b}(X,Y)$.
@@
The inclusion ${\cal L}_{otb}(X,Y)\subseteq{\cal L}_{b}(X,Y)$ follows from [15, Corollary 2.9]. The reverse inclusion hods due to normality of $X$. ∎
Another case of automatic continuity of operators concerns their order continuity. Namely, as every order continuous operator from an Archimedean OVS with generating cone to an OVS is order bounded [15, Corollary 2.2], we have: ${\cal L}_{oc}(X,Y)\subseteq{\cal L}_{b}(X,Y)$, whenever $X$ and $Y$ are OBSs with generating cones such that $X_{+}$ is closed and $Y_{+}$ is normal. Condition $X_{+}-X_{+}=X$ is essential here since each $T:X\to X$ is order bounded and order continuous whenever $X_{+}=\{0\}$.
Example 3.10.
([15, Example 2.12 b)]) Choose a normalized Hamel basis $\{\xi:\xi\in\Xi\}$ of an infinite dimensional Banach space $X$ and identify $X$ with $c_{00}(\Xi)$. Consider a positive cone $X_{+}:=\Big\{\sum\limits_{\xi\in\Xi}\lambda_{\xi}\xi:\lambda_{\xi}\geq 0\Big\}$ in $X$. The cone $X_{+}$ is generating. However $X_{+}$ is not closed ((e.g., by [6, Proposition 4.1])). Pick any sequence $(\xi_{n})_{n=0}^{\infty}$ of distinct elements of $\Xi$ and define $T:(X,X_{+})\to X$ by $T\xi=0$ for $\xi\in\Xi\setminus\{\xi_{n}:n\in\mathbb{N}\}$ and $T\xi_{n}=n\xi_{0}$ for $n\in\mathbb{N}$. It is straightforward to see that the operator $T$ is order-to-norm bounded and positive, but $T$ is not bounded.
It is worth nothing that the order-to-norm and order-to-weak continuous operators on vector lattices were defined in [21] without specifying the class of (just linear or continuous linear) operators. Automatic boundedness of such operators was investigated in the recent paper [22].
We are going to establish further conditions for automatic boundedness of order-to-weak continuous operators between ordered Banach spaces and and normed spaces. More precisely, we prove Theorem 3.12 which tells us that every $\sigma$-w-Lebesgue operator from an OBS with a closed generating normal cone to an NS is bounded. Let us begin with conditions under which $\sigma$-w-Lebesgue operators are order-to-norm bounded.
@@
Let $X$ be an ordered Banach space with a normal cone and $Y$ a normed space. Then ${\cal L}^{\sigma}_{wLeb}(X,Y)\subseteq{\cal L}_{otb}(X,Y)$. In particular, ${\cal L}^{\sigma}_{owc}(X,Y)\subseteq{\cal L}_{otb}(X,Y)$.
@@
Let $T:X\to Y$ be $\sigma$-w-Lebesgue. If $T$ is not order-to-norm bounded then $T[0,u]$ is not bounded for some $u\in X_{+}$. Since $X$ is normal then $[0,u]$ is bounded, say $\sup\limits_{x\in[0,u]}\|x\|\leq M$. Pick a sequence $(u_{n})$ in $[0,u]$ with $\|Tu_{n}\|\geq n2^{n}$, and set $y_{n}:=\|\cdot\|$-$\sum\limits_{k=n}^{\infty}2^{-k}u_{k}$ for $n\in\mathbb{N}$. Then $y_{n}\downarrow\geq 0$. Let $0\leq y_{0}\leq y_{n}$ for all $n\in\mathbb{N}$. Since $0\leq y_{0}\leq 2^{1-n}u$ it follows that $\|y_{0}\|\leq M2^{1-n}$ for every $n$. Thus, $y_{n}\downarrow 0$. It follows from $T\in{\cal L}^{\sigma}_{wLeb}(X,Y)$ that $(Ty_{n})$ is w-null, and hence is norm bounded, that is absurd because $\|Ty_{n+1}-Ty_{n}\|=\|T(2^{-n}u_{n})\|\geq n$ for all $n\in\mathbb{N}$. ∎
Combining Lemmas 3.8, 3.9, and 3.11, we obtain.
@@
Let $X$ be an ordered Banach space with a closed generating normal cone and $Y$ a normed space. Then ${\cal L}^{\sigma}_{wLeb}(X,Y)\subseteq{\cal L}_{b}(X,Y)$. In particular, ${\cal L}^{\sigma}_{ows}(X,Y)\subseteq{\cal L}_{b}(X,Y)$.
@@
(cf. [5, Proposition 3.2], [22, Lemma 2.3]) Each $\sigma$-w-Lebesgue, and hence each $\sigma$-order-to-weak continuous, operator from a Banach lattice to an normed space is bounded.
Positive order-to-norm continuous order-to-norm bounded operators between normed lattices are not necessary bounded [5, Remark 1.1. c)]. From the other hand, even a positive $\sigma$-Lebesgue rank one contraction on an AM-space need not to be w-Lebesgue [5, Example 3.1].
@@
The norm in an ordered normed space $X$ is order continuous if $\|x_{\alpha}\|\to 0$ for every net $x_{\alpha}{\xrightarrow[]{o}}0$ in $X$.
@@
Let $X$ be an OBS with a closed normal cone and the norm in $X$ is order continuous. Then, $x_{\alpha}{\xrightarrow[]{o}}0\Longleftrightarrow x_{\alpha}{\xrightarrow[]{ru}}0$ in $X$.